OpenStudy (anonymous):

Each exterior angle of a regular polygon is approximately 51.43. How many sides does the polygon have?

OpenStudy (anonymous):

i think it is S= (n-2)180

OpenStudy (anonymous):

remember you are looking for the number of sides.

OpenStudy (anonymous):

for "N" sided polygon, there are "N" number of exterior angles. sum of all exterior angles should be 360

OpenStudy (freethinker):

to find 1 angle of a regular convex polygon of n sides = 360/n

OpenStudy (anonymous):

so, the general formula for exterior angle for "N" sides would be, \[\theta={360^\circ\over N}\]

OpenStudy (anonymous):

\[51.43=\frac{ 360 }{ n}\]

OpenStudy (freethinker):

yes!

OpenStudy (freethinker):

now do the rest

OpenStudy (anonymous):

this is true for a "Regular convex polygon" only

OpenStudy (anonymous):

n should equal the amount of sides a polygon has right???

OpenStudy (anonymous):

yep. and that is what you are finding here..

OpenStudy (anonymous):

how many sides did you find out?

OpenStudy (anonymous):

im strying to figure it out

OpenStudy (anonymous):

\[51.43={360\over n}\\ n\times51.43=\cancel{n}\times{360\over \cancel{n}} \]

OpenStudy (anonymous):

similarly, divide both sides by "51.43"

OpenStudy (anonymous):

so i divide 360 by 51.43???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i got 6.99 is that correct ???

OpenStudy (anonymous):

now, can you have 6.99 number of sides? NO so, round it to the nearest integer. that means, you have "7" sides to the polygon

OpenStudy (anonymous):

ohh ok i was going to do that thank you :D

OpenStudy (anonymous):

good. yw